# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

### v1

@@ 5.4. Ablation Study
As mentioned earlier, SmartHalo is composed of two components including static analysis and large language models. To answer RQ3, we compare the Dedaub decompiler, LLM (GPT-3.5) and SmartHalo in terms of method boundaries and variable types, to evaluate the contribution of two components to the overall effectiveness of SmartHalo. Note that we do not compare them in terms of contract attributes because LLM is incapable of outputting the contract attribute without our inference prompt (e.g., contract attribute candidate). To conduct this comparison, we further ran the LLM over the evaluation dataset. As can be seen in Table 5, compared with Dedaub decompiler, LLM presents a limited improvement in precision for method boundary identification but improves the precision of variable type inference by 10.01%. Note that the total number of variable types descends rapidly, because state variables and their types are unexpectedly deleted in the optimized code generated by LLM. We manually inspect the optimization failures of LLM. For method boundary identification, although LLM is capable of recovering a part of inherited methods, it also produces incorrect method boundary prediction which undermines the correct method boundaries in the original decompiler output due to the absence of static knowledge prompts (e.g., new variable declarations and return statements). Benefiting from the inference prompts generated by static analysis, SmartHalo improves the precision of method boundaries by 29.23% and enhances the precision of variable type by 12.52%, compared to LLM alone. In conclusion, both static analysis and the large language model significantly contribute to the improvement of decompiler output optimization in SmartHalo.
Table 5. Comparison results between SmartHalo and LLM.
Method border
Variable type
Original decompiler output (Baseline) 50.48% (316/626) 62.25% (676/1086)
LLM without static inference prompt
51.28% (321/626) 72.26% (375/519)
SmartHalo 80.51% (504/626) 84.78% (1349/1601)
5.5. User study
In this subsection, we highlight the practical effectiveness that SmartHalo helps reverse engineers comprehend the contract bytecode. We invite 13 domain experts and divide them into the professional group (i.e., 8 domain researchers with three years of experience) and intermediate group (i.e., 5 domain researchers with two years of experience), for evaluate the practical value of SmartHalo for different levels of reverse engineers. We randomly select 5 pairs of original decompiler output and optimized code of SmartHalo. Further, we invite each expert to review the code pairs and answer the questions listed in Table 6.
We calculated the average scores and variances for Q1-Q6 across the two groups, with the final results presented in Fig 8. The average scores range from 7.63 to 9.60 between the two groups, and both two groups exhibit relatively small variances within a range below 8. Given that a score of 5 represents a neutral attitude, this indicates that participants at different expertise levels demonstrated a relatively consistent positive attitude towards SmartHalo, particularly regarding method identification, type recovery, contract attributes, downstream tasks, program comprehension, and recompilation. Notably, the average scores from the professional group are lower than those from the intermediate group. To understand this, we interview the participants. The intermediate group views SmartHalo’s decompiler output optimization as comparable to that of a real-world domain expert, resulting in a more favorable response from them. However, due to the higher level of expertise, the professional group found the insights provided by SmartHalo to be less informative compared to the intermediate group. Additionally, the variances in the professional group were more stable than those in the intermediate group, which indicates that more experienced participants demonstrated more consistent and reliable recognition degree. Furthermore, we generated box plots to illustrate the score distribution in both groups, as shown in Fig.8. As can be seen, the scores of professional group are more well-distributed than those of intermediate group.
Table 6. The questions list in the user study.
Q1. Do you think the optimized code is more precise in terms of method borders?
Q2. Do you think the optimized code is more accurate in terms of variable types?
Q3. Do you think the contract attribute of optimized code is meaningful?
Q4. Do you think the optimized code is more helpful in comprehending bytecode?
Q5. Do you think the optimized code help the downstream program analysis tasks?
Q6. Do you think recompilation is an important improvement for decompiled code?
Figure 7. The average scores and variances.
Figure 8. The boxplot for the scores across two groups.
6. Effectiveness for Downstream Task
To evaluate the effectiveness of SmartHalo in supporting downstream program analysis tasks, we conducted three essential experiments. These experiments were designed to assess how effective the contract attributes, method boundaries, and variable types recovered by SmartHalo contribute to their respective downstream tasks.
Contract attributes for Reentrancy vulnerability detection.  As discussed earlier in Section 2.2, contract attribute is critical for locating certain types of vulnerability (e.g, Reentrancy). Prior research (Liao et al., 2022) shows that the effectiveness of contract attribute recovery reflects on eliminating the false positives for Reentrancy detection. Hence, We evaluate how the contract attributes optimized by SmartHalo contribute to improving the precision of Reentrancy detection. To this end, we utilize the latest manually-labeled DApp dataset introduced in previous research (Zheng et al., 2024), with a total of 81 positive labels for Reentrancy vulnerabilities. To conduct this evaluation, we run the SOTA tool SliSE (Wang et al., 2024) and another tool that integrates SliSE with SmartHalo over the manually-labeled DApp dataset to evaluate their precision. As can be seen in Table 6, SliSE+SmartHalo improve the precision to 80.41%, whereas the precision of SliSE is only 72.16%. We further manually inspect all the false positives reported by these two tools. We found that 8 of 27 false positives reported by SliSE can be eliminated by SmartHalo, with the help of more accurate contract attributes recovered by SliSE+SmartHalo. Fig. 9(a) illustrates a Reentrancy example of a false positive reported by SliSE. In this instance, SliSE erroneously flags the function recover due to its adherence to a pre-defined Reentrancy pattern, where function recover modifies state variables after transferring the asset (lines 8 and 9). Upon meticulous inspection of the source code, we found that the contract attribute of the state variable RecoveryDisable belong to [non-asset] (line 02). This attribute indicates that the recover cannot be exploited for profit gain by adversaries, thus recover does not pose a Reentrancy risk. In contrast, SmartHalo accurately identifies the contract attribute of this state variable as [Flag] (line 10), effectively eliminating this false positive.
@@ 6. Effectiveness for Downstream Task
Table 7. Comparison results for evaluating the effectiveness of SmartHalo on the Reentrancy detection
@@
Table 8. Comparison results for evaluating the effectiveness of SmartHalo on the attack identification.
@@
Table 9. Comparison results for evaluating the effectiveness of SmartHalo on Integer-overflow vulnerability.
@@
Method identification for attack identification.  Prior research (Yang et al., 2024) shows that accurate method identification is critical for locating certain types of tasks, such as attack identification. Specifically, attack identification relies heavily on the precise cross-contract call flow analysis between victim and attacker contracts, which effectively models the attack path. Incorrect method identification can lead to missed external calls and corresponding call flows, thereby impeding the identification of attack path. Therefore, the effectiveness of method boundary optimization is reflected in reducing false negatives in attack identification. To assess how method boundaries optimized by SmartHalo enhance the recall of attack identification, we utilize the latest attack contract datasets introduced by previous research (Yang et al., 2024), comprising 18 pairs of attacker and victim contracts from 15 real-world incidents. Further, we compare the SOTA tool BlockWatchdog (Yang et al., 2024) with an enhanced version that integrates SmartHalo with BlockWatchdog. As shown in Table 6, BlockWatchdog+SmartHalo enhances the recall by 16.67% compared with BlockWatchdog alone. Additionally, we manually inspect all the false negatives reported by BlockWatchdog, finding that BlockWatchdog+SmartHalo successfully eliminates all false positives through more accurate method identification optimized by SmartHalo. Fig. 9 (b) presents an example of false negative missed by BlockWatchdog. In this case, function LCtimeout contains a Reentrancy vulnerability, where it logs the transfer results (line 4) after transferring the asset (line 3). Further, an adversary could exploit this by invoking the function LCTimeout to trigger the Reentrancy vulnerability (path $A_{p1}$). BlockWatchdog fails to detect this attack path (i.e., false negative), because function LCtimeout is unavailable for BlockWatchdog at the decompiled code level (Yang et al., 2024). In contrast, SmartHalo accurately identify the method LCTimeout (line 13-15), thereby eliminating this false negative.
Variable types for Integer-overflow vulnerability detection.  Prior works demonstrate that variable type is crucial for locating certain types of overflow vulnerability, such as Integer overflow (Lai and Luo, 2020). For instance, detecting integer-overflow vulnerabilities heavily depends on accurately identifying the upper and lower bounds of variables based on their types. Incorrect variable type recovery results in both false positives and false negatives  (Lai and Luo, 2020). Therefore, the effectiveness of variable type optimization is reflected in its ability to reduce false negatives and false positives in integer-overflow vulnerability detection. We further evaluate the effectiveness of the variable types optimized by SmartHalo on the precision and recall of integer-overflow vulnerability detection. For this evaluation, we utilize integer overflow datasets collected from real-world attacks in previous research (Zhou et al., 2020), which includes 50 vulnerable contracts with integer overflow vulnerabilities. Further, we compare the SOTA tool Mythril (Consensys, [n. d.]) with a modified version that integrates Mythril with SmartHalo. As shown in Table 6, Mythril+SmartHalo enhances the precision by 21.96% and the recall by 38.00% compared with Mythril alone. Additionally, we manually investigate all the false negatives and false positives reported by Mythril. We found that 60% of the false positives and 51.35% of the false negatives reported by Mythril can be eliminated by Mythril+SmartHalo, thanks to the more accurate variable types optimized by SmartHalo. Fig. 9 shows an example of false negative missed by Mythril. In this case, the function batchTransfer contains a uint256 variable amount, which is correlated with two arguments, receivers and value, both of which can be passed arbitrarily via external calls (line 4), leading to an integer overflow vulnerability. Mythril fails to identify this vulnerability on the decompiled code level (i.e., false negative), as the decompiled code does not provide the type information for variable v3 (line 08). Conversely, SmartHalo correctly identifies the type of variable v3 as a uint256 type (line 7), thus eliminating the false negative.
@@
Figure 9. The false alarms reported by SOTA tools, which can be actually eliminated by modified versions that integrate SOTA tools with SmartHalo.
7. Discussion and Limitation

### v2

@@ V-D Ablation Study
As can be seen in Table V, compared with Dedaub decompiler, LLM presents a limited improvement in precision and recall for function boundary identification. While LLM improves the precision of variable type inference by 10.01%, but LLM drop the recall of variable type inference by 18.80%. We manually inspect the optimization failures of LLM. Our manual investigation results shows that, (1) for function boundary identification, although LLM is capable of recovering a part of inherited functions, it also produces incorrect function boundary prediction which undermines the correct function boundaries in the original decompiler output due to the absence of static knowledge prompts. (2) for variable type inference, the recall of SmartHalo decreases rapidly, because a certain number of state variables and their types are unexpectedly deleted in the optimized code generated by LLM alone.
TABLE VI: Comparison results over complex contract dataset, for evaluating SmartHalo’s feasibility to real-world contracts.
Metrics Function boundary Variable type Contract attribute
TP FP FN Prec. Recall TP FP FN Prec. Recall TP FP FN Prec. Recall
Original decompiler
519 223 381 69.95% 57.67% 2955 867 874 77.31% 76.18% N/A
SmartHalo 703 267 197 72.47% 78.11% 3607 222 272 94.20% 92.98% 433 65 464 86.95% 48.27%
“N/A” : the corresponding tool does not support optimizing this deficiency.
Benefiting from the inference prompts generated by static analysis, SmartHalo improves the precision and recall of function boundaries by 19.23% and 29.23%, as well as enhances the precision and recall of variable type by 15.01% and 60.84%, compared to LLM alone. In conclusion, static analysis and the large language model collaboratively contribute to the improvement of decompiler output optimization in SmartHalo.
V-E Generalizability Ability
To answer RQ4, we evaluate the generalizability of SmartHalo in terms of three aspects : 1) real-world complex contract, 2) cross-LLM applicability and 3) cross-decompiler adaptability.
Feasibility to complex and diverse contracts.  As discussed in Sections IV-A and IV-B, SmartHalo has two advantages: 1) both its SA and the LLM components are designed to be general, which enable SmartHalo to address diverse and complex real-world scenarios effectively; and (2) the prompts generated by static analysis embed rich code-context information which captures all the contents relevant to the optimization targets across the whole contract, which enables SmartHalo to perform contract-level optimization.
To evaluate this feasibility, we utilize the complex-contract dataset introduced by previous research [37], which aggregates 682 real-world DApps from 1,199 audit reports. From this corpus, we randomly sampled 50 smart contracts (i.e., around 900 functions in total) for evaluation. Manual investigation of these 50 contracts confirmed that (1) they are highly complex, which extensively cover inheritance, modifiers, and complex storage patterns; and (2) they are sufficiently diverse, which possess more than 12 mainstream smart-contract scenarios.
As shown in Table VI, SmartHalo exhibits good performance on all evaluation metrics when optimizing these complex and diverse contracts at whole-contract optimization granularity. For example, SmartHalo improves the recall of function boundary identification by 20.44%; increases the precision and recall of variable type inference by 16.89% and 16.80%; and achieve the recovery of contract attributes with a precision of 86.95% and a recall of 48.27%, whereas the original decompiled outputs lacks the contract attributes. In summary, SmartHalo can be effectively applied to a wide range of complex real-world contract types, and it exhibits good scalability to enables contract-level decompilation optimization.
TABLE VII: The LLMs evaluated in our study.
Model Type Model Name Open-source Reasoning Size
General LLM GPT-3.5 -
GPT-4o mini -
Llama-3 (Local) 7B
Deepseek-v3 671B
Code LLM Qwen-2.5-coder 32B
Applicability to different LLMs.  Here, we further investigate SmartHalo’s generalizability on integrating different LLMs. The study of the kind of LLM that can better drive SmartHalo is not the focus of our work. Conversely, we aim to explore whether SmartHalo maintains a good performance across different LLMs. To this end, we selected a diverse set of LLMs for evaluation, which considers dimensions such as general LLM versus code LLM, open-source versus closed-source models, reasoning-enabled versus non-reasoning models, and different versions of the same LLM (see Table VII). Among these LLMs, “Llama 3” is actually a small-scale model (7B parameters) that we deploy in local environment.
As shown in Table VIII, SmartHalo maintains good performance across the three metrics when applying to different LLMs. For GPT-4o mini, the precision and recall of function boundary identification are improved by 23.26% and 36.90%, the precision and recall of variable type inference are increased by 28.15% and 46.60%, and the contract attributes are successfully identified with a precision of 83.66% and a recall of 90.93%. For Deepseek-R1, the recall of function boundary identification is enhanced by 41.37%, the precision and recall of variable type inference are increased by 19.47% and 28.42%, and the contract attributes are successfully recovered with a precision of 83.91% and a recall of 86.11%. For Qwen-2.5-coder, the precision and recall of function boundary identification are improved by 95.60% and 85.90%, the precision and recall of variable type inference are increased by 23.89% and 13.31%, and the contract attributes are successfully identified with a precision of 70.54% and a recall of 88.88%. In particular, SmartHalo continues to perform well on the locally-deployed 7B-parameter Llama-3 model ( see the third row of Table VIII), which indicates that SmartHalo can support nearly zero-cost local deployment.
In summary, SmartHalo exhibits good generalizability for a wide range of LLMs regardless of their model type, open-source status, reasoning capability, or model version.
TABLE VIII: Evaluation results across SmartHalo’s versions with different LLMs, for evaluating the SmartHalo’s applicability to different LLMs.
Setup Function boundary Variable type Contract attribute
TP FP FN Prec. Recall TP FP FN prec. recall TP FP FN prec. recall
SmartHalo with GPT-3.5 504 67 122 88.26% 80.51% 1349 113 252 92.27% 84.26% 752 350 75 68.06% 90.93%
SmartHalo with GPT-4o mini 547 52 79 91.32% 87.38% 1422 151 179 90.40% 88.82% 759 182 68 80.66% 91.78%
SmartHalo with Llama 345 177 281 66.09% 55.11% 777 468 824 62.41% 48.53% 499 310 328 61.68% 60.34%
SmartHalo with Deepseek-R1 575 397 51 59.16% 91.85% 1131 253 470 81.72% 70.64% 626 120 201 83.91% 86.11%
SmartHalo with Qwen 544 25 82 95.60% 86.90% 889 143 712 86.14% 55.53% 735 307 92 70.54% 88.88%
TABLE IX: Evaluation results across different decompilers, for evaluating SmartHalo’s adaptability to to different decompilers.
Metrics Function boundary Variable type Contract attribute
TP FP FN prec. recall TP FP FN prec. recall TP FP FN prec. recall
Dedaub
316 149 310 67.96% 50.48% 676 410 975 62.25% 42.22% N/A
SmartHalo on Dedaub 504 67 122 88.26% 80.51% 1349 113 252 92.27% 84.26% 752 350 75 68.06% 90.93%
Heimdall
316 149 310 67.96% 50.48% 984 1082 617 47.63% 61.46% N/A
SmartHalo on Heimdall 440 109 186 80.15% 70.28% 1502 834 99 64.30% 93.82% 823 942 4 46.63% 99.52%
Panoramix
225 149 401 60.16% 35.94% 583 348 1018 62.62% 36.41% N/A
SmartHalo on Panoramix 368 88 258 80.70% 58.78% 1281 162 320 88.77% 80.01% 677 151 150 81.76% 81.86%
“N/A” : the corresponding tool does not support optimizing this deficiency.
Adaptability to different decompilers.  Lastly, we evaluate SmartHalo’s ability to deal with discrepancies among different Solidity decompilers. To this end, we selected three mainstream decompilers, Gigahorse (i.e., Elipmoc), Heimdall, and Panoramix, all of which have been widely applied in prior studies and industry [38, 39]. Specifically, we run these three decompilers on the 456 functions contained in our evaluation dataset to obtain their respective decompiled outputs. Subsequently, SmartHalo is run on the decompiled code produced by each decompiler, and the results of three evaluation metrics are collected for comparative analysis.
As illustrated in Table IX, evaluation results show that SmartHalo improves both precision and recall across all the metrics for different decompilers. For Heimdall, the precision and recall of function boundary identification are enhanced by 12.19% and 19.80%, the precision and recall of variable type inference are increased by 16.67% and 32.36%, and the contract attributes are successfully recovered with a precision of 46.63% and a recall of 99.52%, whereas the original decompiled outputs lack the contract attributes. For Panoramix, the precision and recall of function boundary identification are improved by 20.54% and 22.84%, the precision and recall of variable type inference are increased by 26.15% and 43.60%, and the contract attributes are successfully identified with a precision of 81.76% and a recall of 81.86%, whereas the original decompiled outputs lack the contract attributes. These results indicate that SmartHalo is generic to different Solidity decompilers.
V-F Efficiency
To address RQ5, we execute SmartHalo on our evaluation dataset and measure both its runtime and monetary costs.
As shown in Table X, the proposed approach can establishes high-quality optimization with an efficient and cost-effective manner. For example, we compute the average time for each step of SmartHalo: 1.94 s for the dependency-based semantic extraction module, 8.95 s for the LLM-driven semantic enrichment module and 13.10 s for the correctness verification module. For monetary costs, the GPT-4o mini-powered version of SmartHalo consumes an average of 8,920 tokens for each function, which incurs an average expense of \$0.00136. With our statistical analysis, the average number of functions per contract within the evaluation dataset is approximately 34. Accordingly, we can roughly estimate that, SmartHalo costs an average expense of \$0.046 and an average runtime of 815.66 s per smart contract.
In addition, with the rapid development of LLMs, an increasing number of open-source alternative LLMs has emerged. As mentioned earlier, SmartHalo can be deployed with a local, cost-free, small-scale LLama-7B model and achieve good performance, as shown in Table VIII. In this way, SmartHalo’s monetary cost can be reduced to nearly zero.
In sum, SmartHalo is proved to be cost-effective in terms of runtime and monetary costs, which indicate its practical value.
Parameter setting influence.  Further, we evaluate how the settings of iteration limit influence the cost of correctness verification. Specifically, we run the correctness verification module with different iteration limit (n=1,2,3n=1,2,3…), and compute the convergence rate, runtime and monetary costs of evaluated functions. The evaluation results show that, when $n\geq 3$, both the runtime and monetary costs stop increasing, as the convergence rate of correctness verification module reaches 100%. Hence, we set the iteration limit $n$ to 3 in our evaluation setup.
TABLE X: The average time and tokens for analyzing each method in the evaluation dataset.
Efficiency of SmartHalo Avg.time(s) Avg.tokens
Dependency-based semantic extraction 1.94 -
LLM-driven semantic enrichment 8.95 4272 (\$ 0.00065)
Correctness verification 13.10 4648 (\$ 0.00071)
Total 23.99 8920 (\$ 0.00136)
@@ V-F Efficiency
TABLE XI: Comparison results for evaluating the effectiveness of SmartHalo on the Reentrancy detection
@@
TABLE XII: Comparison results for evaluating the effectiveness of SmartHalo on the attack identification.
@@
TABLE XIII: Comparison results for evaluating the effectiveness of SmartHalo on Integer-overflow vulnerability.
@@
VI Effectiveness for Downstream Task
These experiments were designed to assess how the contract attributes, function boundaries, and variable types recovered by SmartHalo contribute to their respective downstream tasks.
Contract attributes for Reentrancy vulnerability detection.  Prior research [5] shows that the effectiveness of contract attribute recovery reflects on eliminating the false positives for Reentrancy detection. Hence, We evaluate how the contract attributes optimized by SmartHalo contribute to improving the precision of Reentrancy detection. To this end, we utilize the latest manually-labeled DApp dataset introduced in previous research [40], with a total of 81 positive labels for Reentrancy vulnerabilities. To conduct this evaluation, we run the SOTA tool SliSE [40] and another tool that integrates SliSE with SmartHalo over the manually-labeled DApp dataset to evaluate their precision. As can be seen in Table V-F, SliSE+SmartHalo improve the precision to 80.41%, whereas the precision of SliSE is only 72.16%. We further manually inspect all the false positives reported by these two tools. We found that 8 of 27 false positives reported by SliSE can be eliminated by SmartHalo, with the help of more accurate contract attributes recovered by SliSE+SmartHalo. Fig. 8(a) illustrates a Reentrancy example of a false positive reported by SliSE. In this instance, SliSE erroneously flags the function recover due to its adherence to a pre-defined Reentrancy pattern, where function recover modifies state variables after the external call (lines 8 and 9). Upon meticulous inspection of the source code, we found that the contract attribute of the state variable RecoveryDisable belong to [non-asset] (line 02). This attribute indicates that the recover cannot be exploited for profit gain by adversaries, thus recover does not pose a Reentrancy risk. In contrast, SmartHalo accurately identifies the contract attribute of this state variable as [Flag] (line 10), effectively eliminating this false positive.
@@ VI Effectiveness for Downstream Task
function identification for attack identification.  Prior research [7] shows that attack identification relies heavily on the precise cross-contract call flow analysis between victim and attacker contracts, which effectively models the attack path. Incorrect function identification can lead to missed external calls and corresponding call flows, thereby impeding the identification of attack path. Therefore, the effectiveness of function boundary optimization is reflected in reducing false negatives in attack identification. To assess how function boundaries optimized by SmartHalo enhance the recall of attack identification, we utilize the latest attack contract datasets introduced by previous research [7], comprising 18 pairs of attacker and victim contracts from 15 real-world incidents. Further, we compare the SOTA tool BlockWatchdog [7] with an enhanced version that integrates SmartHalo with BlockWatchdog. As shown in Table V-F, BlockWatchdog+SmartHalo enhances the recall by 16.67% compared with BlockWatchdog alone. Additionally, we manually inspect all the false negatives reported by BlockWatchdog, finding that BlockWatchdog+SmartHalo successfully eliminates all false positives through more accurate function identification optimized by SmartHalo. Fig. 8 (b) presents an example of false negative missed by BlockWatchdog. In this case, function LCtimeout contains a Reentrancy vulnerability, where it logs the transfer results (line 4) after transferring the asset (line 3). Further, an adversary could exploit this by invoking the function LCTimeout to trigger the Reentrancy vulnerability (path $A_{p1}$). BlockWatchdog fails to detect this attack path (i.e., false negative), because function LCtimeout is unavailable for BlockWatchdog at the decompiled code level [7]. In contrast, SmartHalo accurately identify the function LCTimeout (line 13-15), thereby eliminating this false negative.
@@
Figure 8: The sample false alarms reported by SOTA tools. These FPs and FNs can be eliminated with the integration of SmartHalo.
Variable types for Integer-overflow vulnerability detection.  Prior research [26, 41] demonstrate that detecting integer-overflow vulnerabilities heavily depends on accurately identifying the upper and lower bounds of variables based on their types. Incorrect variable type recovery results in both false positives and false negatives  [26]. Therefore, the effectiveness of variable type optimization is reflected in its ability to reduce false negatives and false positives in integer-overflow vulnerability detection. We further evaluate the effectiveness of the variable types optimized by SmartHalo on the precision and recall of integer-overflow vulnerability detection. For this evaluation, we utilize integer overflow datasets collected from real-world attacks in previous research [42], which includes 50 vulnerable contracts with integer overflow vulnerabilities. Further, we compare the SOTA tool Mythril [43] with a modified version that integrates Mythril with SmartHalo. As shown in Table V-F, Mythril+SmartHalo enhances the precision by 21.96% and the recall by 38.00% compared with Mythril alone. Additionally, we manually investigate all the false negatives and false positives reported by Mythril. We found that 60% of the false positives and 51.35% of the false negatives reported by Mythril can be eliminated by Mythril+SmartHalo, thanks to the more accurate variable types optimized by SmartHalo. Fig. 8 shows an example of false negative missed by Mythril. In this case, the function batchTransfer contains a uint256 variable amount, which is correlated with two arguments, receivers and value, both of which can be passed arbitrarily via external calls (line 4), leading to an integer overflow vulnerability.
Mythril detects integer overflow based on two conditions: (1) scanning for key EVM instructions (e.g., ADD, SUB, MUL, etc.), and (2) invoking Z3 solver with the maximum bit-width derived from the variable type to determine whether integer overflow occurs. Due to the absence of type information for the variable v3 (line 07) in the decompiled code, Mythril fails to identify the maximum bit-width, resulting in a false negative report. Conversely, SmartHalo correctly identifies the type of variable v3 as a uint256 type (line 07), thus eliminating the false negative.
VII Discussion and Limitation